OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4)

OpenStudy (help!!!!):

use point-slope

OpenStudy (anonymous):

paralell linnes have same slope so resultant line will also have slope same as given line.. so that is m=slope=4 so now use this.. (y-y1)=m(x-x1) plug in values ..

OpenStudy (help!!!!):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (help!!!!):

What I would first do is change the standard form to y-intercept. Then use point-slope form for the 2 coordinates.

OpenStudy (anonymous):

i did all that and got a answer that isnt one of the options

OpenStudy (anonymous):

y = four thirdsx + 8 y = four thirdsx + 3 y = -three fourthsx + 8 y = -three fourthsx + 3

OpenStudy (anonymous):

Graph it and translate the line

OpenStudy (help!!!!):

Its the first one

OpenStudy (anonymous):

-3y-4=4(x+3) -3y-4=4x+12 -3y=4x+16 y=-4/3x+16/3

OpenStudy (help!!!!):

your slope you wrote down is incorrect, but rather 4/3

OpenStudy (help!!!!):

redo the standard form to y-intercept form

OpenStudy (anonymous):

y = 4/3x + 8

OpenStudy (help!!!!):

exactly the first choice

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

okay thanks

OpenStudy (help!!!!):

np